light down the changes you would observe if speed of light is 1m/s and the changes if you move with half the speed of light
1
Expert's answer
2014-06-07T05:12:10-0400
Change of fundamental constants like speed of light c, the Plank constant h or electron charge e results in the change of the fine-structure constant alpha (https://en.wikipedia.org/wiki/Fine-structure_constant). It can be interpreted in many ways in physics. Considering this question, it's important that alpha characterizes the strength of the electromagnetic interaction. Strong change of the fine-structure constant makes impossible stability of atoms and nuclear reactions in stars. Decreasing of the speed of light from 300,000 km/s to 1 m/s increases alpha. According to the interpretations of the quantum electrodynamics, it means that the characteristic energy of interaction also become larger. It is one of ways to consider process of merging of all known interactions (electromagnetic, strong, weak and gravitational) into the one for time of Universe birth.

When some object moves with speed comparable with the speed of light, it results in the relativistic reduction of the object's length and it's own time according to the Lorentz transformations (https://en.wikipedia.org/wiki/Lorentz_transformation). In other words, the clocks of the object will show time smaller in \sqrt{1-v^2/c^2} = \sqrt{1 - 0.5^2} = 0.87 times than stationary clocks.